You push a shopping cart 200.0 m by exerting a force of 12.00 N onto the handle at an angle of 20o above the horizontal axis. If the cart experiences a force of friction of 2.00 N, calculate the total amount of work done on the cart.


1
Expert's answer
2021-03-12T07:13:54-0500

Let's find the work done by the pushing force:


"W_p=F_pdcos\theta=12.0\ N\cdot200.0\ m\cdot cos20^{\circ}=2255\ J."

Let's find the work done by the force of friction:


"W_{fr}=F_{fr}dcos\theta=2.0\ N\cdot200.0\ m\cdot cos180^{\circ}=-400\ J."

Finally, we can find the net (total) work done on the cart:


"W_{net}=W_p-W_{fr}=2255\ J-400\ J=1855\ J."